OpenStudy (anonymous):

What is the solution to the equation? 4(3 + 7) = 4 · x + 4 · 7

OpenStudy (anonymous):

1)4*3 and 4*7 2)4*1x and 4*7 3)12+28=4x+28 4)40=4x+28 5) Subtract 28 on both sides 40-28 is 12 6)4x=12 7)Divide by 4 8)Answer is 3